OpenStudy (anonymous):

For how many integer values of x^2+ax+6 a can be factored? What are they?

OpenStudy (anonymous):

i need help on this

OpenStudy (anonymous):

\[\mathrm{x^2+ax+6=(x+m)(x+n) \\x^2+ax+6=x^2+(m+n)x+mn}\] \[\mathrm{m+n=a\\ mn=6}\] I hope this helps.

OpenStudy (anonymous):

it does alittle

OpenStudy (anonymous):

xD

OpenStudy (anonymous):

\[\mathrm{m(a-m)=6=2\times3=6\times1}\]

OpenStudy (anonymous):

You have to consider the negative integers :p

OpenStudy (anonymous):

so is the awnser -7 -5 5 7

OpenStudy (anonymous):

4 integer values; -7, -5, 5, and 7

OpenStudy (anonymous):

Yes :p

OpenStudy (anonymous):

i dont know what the work looks like

OpenStudy (anonymous):

anyone there

OpenStudy (anonymous):

X^2 + ax + 6 can be factored either in the form (x - b)(x - c) or (x + b)(x + c). b and c can any pair of the factors of 6: 4 integer values; -7, -5, 5, and 7 i hope this help..

OpenStudy (anonymous):

\[m(a-m)=6\\ a=\frac{6}{m}+m,\quad m\neq0\\ \] Integer+rational=rational If m is integer, then 6/m is integer (if m=2,3,6) or rational. If m is rational, then 6/m is integer. Then a is rational hence these are all the solutions (what you said before)